Tohban Report 2010-05-26

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 19 May 2010
End Date: 26 May 2010
Tohban: Pascal Saint-Hilaire
Tohban email:
Next Tohban: Pascal Saint-Hilaire
List all reports

Solar Activity

GOES level steadily increased from A2 to A4 level. A few B-flares occurred.

How many GOES flares occurred?

 Flares above B, C, M, X class were      5     0     0     0

Memory Management

Despite a relatively quiet Sun, we are having problems emptying the SSR. This week, SSR varied from 41 to 76%. A 12-hour chunk (doy 140) and a 7.5-hour (doy 143) chunks of data were deleted, in order to bring the SSR to a lower level. It appears that 18 fully healthy segments are the main cause. The previously-reported light leak in the back of cryostat turned out to be not such a big issue after all, and seem to be diminishing with S/C precession. Detector 5 had caused some concern last week or so, but is now very well behaved.

We will have a few more Waldheim passes this week (2/day, 10 total), but we cannot count on getting much more than 9-10 passes per day on a long term basis. David Smith is going to try re-trimming SAA boundaries, in order to collect less SAA counts.

For the time being, it has been decided to shut off rear detectors at night. It is expected we could gain as much as ~20% memory per orbit. We remain at Normal/Vigorous decimation schemes, fronts off during S/C night, and both magnetic zones are heavily decimated.

The goal is too have a very low SSR fill level when Crab operations start on June 7th.

Data Gaps

Quite a few of them, actually (besides the afore-mentioned data deletions). They look to be all Wallops issue, but data should be on the ground. Jeremy is going to look into this.

Personal tools